A regression was run to determine if there is a relationship between hours of TV watched per day (x) and number of situps a person can do (y).

The results of the regression were:

y=ax+b
a=-1.005
b=33.926
r2=0.3969
r=-0.63
n=30



Use this to predict the number of situps a person who watches 6.5 hours of TV can do (to one decimal place)
Answer #1

Since given regression equation is y = ax+b where a= -1.005 and b = 33.926

So, y = -1.005*X + 33.926 where y is number of situps a person can do and X is hours of TV watched per day.

So when X = 6.5 then Y = -1.005*6.5 + 33.926 = 27.4

So the number of situps a person who watches 6.5 hours of TV can do is 27.4